Please walk me through the steps to the following problem. I don't even know how to start.
A Ford Explorer can hold a driver and four passengers. A Dodge Caravan can hold a driver and six passengers. The biology dept. is taking a total of 95 people on a field trip. The college wants to use twice as many Caravans as Explorers because the Caravans get better gas mileage. How many of each kind of vehicle are needed?

Answer by mananth(16949)

A Ford Explorer can hold a driver and four passengers. A Dodge Caravan can hold a driver and six passengers. The biology dept. is taking a total of 95 people on a field trip. The college wants to use twice as many Caravans as Explorers because the Caravans get better gas mileage. How many of each kind of vehicle are needed?
let explorers be x
and caravans be y
y=2x
5x+7y=95
substitute y=2x
5x+7(2x)=95
5x+14x=95
19x=95
x=5
5 explorers and 10 caravans